# Redistricting Reforms Reduce Gerrymandering by Constraining Partisan Actors

### v1

@@ 3.1 Data on the relevant institutional features
Our standardized coding of relevant institutional features is based on the prototypical redistricting process shown in Figure 2. First, an initial map drawer proposes a plan, which may be vetoed by other actors. If the plan is not vetoed, it can be challenged in court or not. If it is vetoed, there is another round of map drawing. If a plan is vetoed twice, or if the initial map drawer cannot agree on a plan, then a different institution (often a court) must resolve the stalemate and adopt a plan.
Every state’s process can be described as a subset of this prototype. For example, in Alabama, the legislature draws congressional districts, subject to a governor’s veto, and there is no mechanism for state court review. Thus, Alabama’s process would be described by the “Drawer” and “Veto 1” steps only in Round 1 of the Figure 2 process.
For each state, we record which institutional body, if any, acts at each step, and which party, if any, controls that institution. We also collected additional information relevant to modeling redistricting processes and court review, such as whether the state’s redistricting plans were subject to DOJ preclearance before 2013,333The Shelby decision was handed down in 2013, ending DOJ preclearance. and which institutional body ended up drawing the plan that was used in the first postcensal elections. Most procedural details are straightforward, and the information for coding is readily available in public data from each state.444For instance, from https://redistricting.lls.edu/. Furthermore, many states have similar redistricting procedures that are easily classified under the categories we defined earlier.
@@ 3.2 The redistricting game
Coding the details of each state’s redistricting process preserves important procedural information, compared to simply categorizing each state into a small number of groups such as “legislature-controlled” and “independent commission.” However, the detailed coding presents a challenge for causal inference, since the treatment—a state’s redistricting process—is now high-dimensional. Out of the 87 state-decade processes we code,555Only 43 states had two or more congressional districts in the 2010 cycle; for 2020, Montana did as well, bringing the total to 44. there are 58 distinct combinations of procedural variables, 41 of which are completely unique.
@@
The combination of high-dimensional treatment and limited sample size means that there may not be enough information to estimate the causal effect of changing from one specific configuration of procedural variables to another, without further assumptions.
@@
We address this methodological challenge by leveraging two basic substantive assumptions about the redistricting process: first, that each party aims to draw a map that favors it as much as possible, and second, that the parties are constrained by statutory and constitutional rules in doing so. Specifically, we treat the redistricting process depicted in Figure 2 as a sequential zero-sum game666The zero-sum assumption rules out cooperation between the parties to, e.g., protect incumbents. Given the role of the game here as a data summary rather than predictive model, we believe this is an acceptable limitation. with two players—the Democratic and Republican parties—each trying to maximize the degree to which the drawn plan favors their party.
@@
In each state, the nodes in the game tree can be controlled by different parties, or by neither party, e.g., when a supermajority is required to adopt a plan and neither party has supermajority control. Any split-control nodes, as well as a node for state court action, are considered moves by nature. Some nodes involve discrete choices such as whether to veto a plan or not, others are labeled “choose bias,” meaning that the player at that node draws a plan with a chosen amount of partisan bias favoring either party. This bias is exactly the utility to the party of the chosen plan: a plan with bias is worth to the Republicans and to the Democrats. We need not quantify exactly what the utility measure is as a function of a specific plan chosen; it suffices to let the parties try to maximize an abstract univariate measure of partisan bias. We let the bias score range from , indicating a maximum Democratic advantage, to , indicating a maximum Republican advantage.
@@
For concreteness, consider the case of redistricting in Alabama, as described earlier. The first move belongs to the Republican party, which controls the state legislature. It has to pick the amount of partisan bias in the plan that it adopts. If this plan is ultimately the final enacted plan, the Republicans receive utility and the Democrats receive utility . The second move also belongs to the Republican party, which controls the governorship. The party must decide whether or not to veto the adopted plan; if it does, there is a second round of drawing and vetoing by the legislature and governor (controlled by Republicans).
@@
All in all, the game specification depends on 19 parameters which govern the moves by nature, with most of these parameters relating to the court challenge process. Rather than fix these parameters to arbitrary constants, we place a prior distribution on each parameter over a range of probable values. We simulate 100 different draws from this joint prior distribution; each draw generates a slightly different game specification. We then average our results across the random draws.
@@ 3.3 Equilibrium solution as a treatment variable
To solve the problem of a high-dimensional treatment, we use the subgame perfect Nash equilibrium of the redistricting game. The utility of the equilibrium solution captures the expected partisan bias of plans that arise out of a state’s redistricting process, under the current party control of the state’s institutions. All of the multi-step institutional interactions and negotiations that might happen as part of the redistricting process are thus reduced to a univariate score. The upshot is that the 14 procedural variables can be reduced into a single axis that measures the leeway political actors have over the partisan lean of the final redistricting plan.
The expected outcome of both the stalemate-then-court and court-alone choices are determined by the parameters governing the moves by nature. They are shown as a function of the bias of the legislature’s proposed plan in Figure 3(a), for a typical set of game parameters. In this case, because the only realistic avenue for court review in Alabama is a VRA claim, the expected outcome of a court challenge given an enacted plan is roughly . The expected bias of a plan drawn through stalemate procedures under a Republican court is roughly 0.5, though it also increases monotonically in . Since the governor wishes to maximize the signed partisan bias, she will veto the plan if its bias is less than about 0.5, and not veto it otherwise.
Having solved the Round 2, Veto 1 subgame, we move backwards to the Round 2 drawing subgame. The Republican legislature must decide on the partisan bias of the plan it proposes, or can choose to stalemate by failing to adopt a plan. If it stalemates, then the expected outcome will be given by the green “Veto” line in Figure 3(a), and will depend on the partisan bias of the plan originally adopted by the legislature in Round 1 before being vetoed. If it does not stalemate, then the expected outcome is determined by the dashed equilibrium line in Figure 3(a). To maximize the partisan bias, the legislature at Round 2 will draw a map with bias of , leading to expected bias (utility) of 2.84 for this set of game parameters.
We can then move backwards again and consider the Round 1, Veto 1 subgame. Figure 3(b) shows the expected outcomes under each choice, just as before. In equilibrium, the governor will veto any plan with bias less than 4; regardless of the governor’s decision, the expected bias is 2.84. Finally, the legislature can make any initial proposal, since the expected outcome is the same regardless. Thus, the overall equilibrium is a bias of 2.84. This solution process is automated across all of the states and is carried out on each of the 100 different draws from the prior on the game’s parameters.
@@
For our causal analysis, our first treatment measure is exactly the game’s Nash equilibrium, averaged across the prior. This measure depends on which parties control each node in the game tree. We also generate a second treatment measure which does not depend on current party control. This is calculated by assigning a single party (here, the Democrats777The game is completely symmetric between parties, with one exception: due to racial polarization in the U.S., challenges to redistricting plans under the VRA are much more likely for Republican-leaning plans, and when these challenges prevail the remedy almost always results in a plan more favorable to Democrats. In keeping with the goal of calculating the maximum leeway in each state, we therefore assign states to Democratic control under the counterfactual scenario here. Assigning to Republican control instead would slightly reduce the calculated leeway for those states previously subject to DOJ preclearance.) to control each node in the game tree that belongs to a partisan actor—legislature, governor, or partisan commission—and then recalculating the average Nash equilibrium. We refer to this equilibrium as the maximum leeway of a state’s redistricting process, since it captures the expected bias under a worst-case partisan outcome where all of the levers of state government are controlled by one party. The first treatment measure we refer to as the realized leeway, since it depends on the realized values of party control for each state institution.
@@
As might be expected, states with independent redistricting commissions, such as Michigan, have a realized and maximum leeway of zero. In 2010, Michigan redistricting was entirely controlled by the legislature and governor, who were both Republican. By 2020, the process was reformed to have a strong independent commission. In the new process, the commission is bound by strict criteria, the process allows for court review, and even the commissioners themselves are selected in part by a lottery system. With the reform, Michigan’s maximum leeway changed from the highest observed value to zero, meaning that partisan actors have no control over redistricting outcomes, on average.
Other states’ scores are similarly interpretable, where introducing a commission (e.g., New York, Virginia, or Colorado) drastically reduces its leeway to match states with similar sets of rules in 2020. States that had intervening litigation to clarify the interpretation of state redistricting rules (e.g., New Mexico, North Carolina, or Pennsylvania) see appropriate changes in leeway, even absent a commission. Further, in our party-signed treatment, states that had a flip in party control of the whole system see very large changes in leeway in the correct direction (e.g. West Virginia). Finally, states with minor changes to the total control of state institutions see minor changes in leeway (e.g., North Carolina).
@@ 3.4 Validation
Building treatment variables through a game-theoretic model provides significant advantages in dimension reduction and interpretability, but it also comes with the risk that the model is misspecified. Although it is difficult to completely verify the validity of each component of the model, we take several steps towards empirically validating the model as a whole, showing that the resulting treatment variable predicts the observed outcomes well.
The model was designed to be flexible enough to capture the actual redistricting processes in each state. Any misspecification is therefore due to how the moves by nature are specified. First, we find that the treatment values are not sensitive to specific parameter values, up to monotonic transformations. Across the 100 random draws from the prior, the average pairwise Spearman correlation between the Nash equilibria for each state is above 0.99.888i.e., for each pair of random draws, we calculate the Spearman correlation between the two vectors of calculated state equilibria, then average these correlations across all pairings. This also gives us confidence that the specific choice of prior is not influencing the results.
Second, as shown in Figure 5, there is a meaningful correlation in the expected direction between the treatment values and measures of partisan advantage. The left panel plots the expected share of seats won by Republicans versus the realized leeway measure; recall that the leeway measure is positive for plans that favor Republicans. The right panel shows the partisan bias (King and Browning,, 1987), measured at the state’s baseline vote share, versus realized leeway. Larger values of partisan bias correspond to plans that systematically favor one party, with positive values favoring Republicans.
For both outcome measures, the purely a priori model predictions of the expected partisan bias do in fact correlate with the actual partisan bias of the plans that come out of each state’s redistricting process. To what extent these correlations can tell us about the causal effects of reforms is of course another matter, and the primary question at issue in this paper.
@@
Table 1 compares the most likely final map-drawer, according to these probabilities, to the actual institution that drew the final redistricting plan. The correspondence is generally excellent, considering the predictions are made based on theory alone. There is a tendency, however, for the model to predict court-drawn maps when they are in fact drawn by legislatures. The overall likelihood of a court intervention is itself a model parameter that is varied across random draws, and as noted above the ranking of the states by leeway is basically unchanged by different parameter values. Therefore, we expect this tendency to over-predict court intervention to at most slightly understate the equilibrium value for states with court review, since a higher likelihood of intervention would tend to pull the equilibrium towards zero.
Finally, it is natural to wonder whether a more accurate model could be obtained by estimating the model parameters from the observed data rather than specifying their prior distributions. That is, one could find the parameter values that maximize the correlation between the treatment and outcome values. We do not take this approach in order to cleanly separate out estimation of causal effects from construction of a treatment variable. This enables us to employ a causal identification strategy that does not assume the correct specification of the game theoretic model. Instead, we use the model to summarize a high-dimensional treatment variable in a theoretically informed way without looking at the outcome variable.999We did attempt to fit the game parameters to the observed data for the 2010 cycle alone via maximum likelihood, but we found that the observed data were only minimally informative about most of the parameter values. This is good news, insofar as it implies that a wide range of parameter values are all compatible with the observed data.
@@ 4 Estimating the Causal Effects of Institutional Leeway
where denotes a generic potential outcome variable for state at time with institutional features .101010For identification, it in fact suffices to assume only that for any with ; however, the stronger individual-level assumption is invariant under transformations and ensures that no additional assumptions are required for estimation. Under this assumption, we can simply write the potential outcomes as for the treatment dose .
Our target estimand is the conditional average treatment effect (CATE) for a change from treatment level to given a set of covariates that are not affected by the treatment. The CATE is defined as,
@@
Our covariates include a range of variables that existing work in the redistricting literature identifies as predicting redistricting outcomes and are not affected by the redistricting reforms. We include (pre-treatment) 2008 Democratic presidential vote share, an indicator for being in the South,111111Southern states are defined as Alabama, Arkansas, Florida, Georgia, Kentucky, Louisiana, Mississippi, North Carolina, South Carolina, Tennessee, Texas, Virginia, and West Virginia. This set corresponds to Confederate states, plus Kentucky and West Virginia, which also had a history of Democratic dominance in the 20th century and notable differences between party preferences at the state and national levels. the logarithm of the number of districts in 2020, the change in the number of districts between 2010 and 2020, the logarithm of the average number of state corruption convictions by year between 2000 and 2010, and an indicator for whether states allow ballot initiatives.
Formally, we require the following conditional parallel trend assumption for all :
@@
To further increase the credibility of the conditional parallel trend assumption above, we adjust for the changes in political geography between the two redistricting cycles. Partisan preferences and the geographic distribution of voters can change in different ways within each state over time. For example, Michigan, which experienced reform, became more Republican during our sample period, while Georgia, which did not, became more Democratic. Without this adjustment, an estimated effect of the Michigan reform would be biased by the differential change in the states’ political geographies.

### v2

@@ 3.1 Data on the relevant institutional features
Our standardized coding of relevant institutional features is based on the prototypical redistricting process shown in Figure 2. First, an initial map drawer proposes a plan that may be vetoed by other actors. If the plan is not vetoed, it can be challenged in court or not. If it is vetoed, there is another round of map drawing. If a plan is vetoed twice, or if the initial map drawer cannot agree on a plan, then a different institution (often a court) must resolve the stalemate and adopt a plan.
Every state’s process can be described as a subset of this prototype. For example, in Michigan, a commission draws congressional districts, resolves any stalemates, and there is an explicit mechanism for state court review. Thus, Michigan’s process would be described by the “Drawer” step in Round 1 and “Stalemates” steps only in Figure 2.
For each state, we record which institutional body, if any, acts at each step and which party, if any, controls that institution. We also collected additional information relevant to modeling redistricting processes and court review, such as whether the state’s redistricting plans were subject to DOJ preclearance before 2013,333The Shelby decision was handed down in 2013, ending DOJ preclearance. and which institutional body ended up drawing the plan that was used in the first postcensal elections. Most procedural details are straightforward, and the information for coding is readily available in public data from each state (see e.g., https://redistricting.lls.edu/). Furthermore, many states have similar redistricting procedures that are easily classified under the categories we defined earlier.
@@ 3.2 The redistricting game
Coding the details of each state’s redistricting process preserves important procedural information, compared to categorizing each state into a small number of groups such as “legislature-controlled” and “independent commission.” However, the detailed coding presents a challenge for causal inference, since the treatment—a state’s redistricting process—is now high-dimensional. Out of the 87 state-decade processes we code,444Only 43 states had two or more congressional districts in the 2010 cycle; for 2020, Montana did as well, bringing the total to 44. there are 58 distinct combinations of procedural variables, 41 of which are completely unique.
@@
The combination of high-dimensional treatment and a limited sample size means that there may not be enough information to estimate the causal effect of changing from one specific configuration of procedural variables to another, without further assumptions.
@@
We address this methodological challenge by leveraging two basic substantive assumptions about the redistricting process. First, each party aims to draw a map that favors it as much as possible, and second, the parties are constrained by statutory and constitutional rules in doing so. Specifically, we treat the redistricting process depicted in Figure 2 as a sequential zero-sum game555The zero-sum assumption rules out cooperation between the parties to, e.g., protect incumbents. Given the role of the game here as a summary of institutional features rather than a structural model that predicts the outcome, we believe this is an acceptable limitation. with two players—the Democratic and Republican parties—each trying to maximize the degree to which the drawn plan favors their party.
@@
In each state, the nodes in the game tree can be controlled by different parties, or by neither party (e.g., when a supermajority is required to adopt a plan and neither party has supermajority control). Any split-control nodes, as well as a node for state court action, are considered moves by nature. Moves at one node do not affect the set of actions available at other nodes. Some nodes involve discrete choices such as whether to veto a plan or not, others are labeled “choose bias,” meaning that the player at that node draws a plan with a chosen amount of partisan bias favoring either party. This bias is exactly the utility for the party of the chosen plan: a plan with bias is worth to the Republicans and to the Democrats. We need not quantify exactly what the utility measure is as a function of a specific plan chosen; it suffices to let the parties try to maximize an abstract univariate measure of partisan bias. We let the bias score range from , indicating a maximum Democratic advantage, to , indicating a maximum Republican advantage.
@@
For concreteness, consider the case of redistricting in Michigan, as described earlier. The first move belongs to the independent commission. It has to pick the amount of partisan bias in the plan that it adopts. If this plan is ultimately the final enacted plan, Republicans receive utility and Democrats receive utility . The stalemate move also belongs to the commission.
If a plan is adopted in the first round, it proceeds to a possible court review. The court may decide to accept a legal challenge, decide in favor of the plaintiffs, and redraw the map; this choice is considered a move by nature. If the court review results in a redrawn map with partisan bias , then the Republicans receive utility and the Democrats receive utility . In Michigan, court review is explicitly allowed on partisan grounds and challenges under the federal Voting Rights Act are possible, so there is moderate probability that the commission-adopted plan will be overturned. If no plan is enacted by the commission in the first round, the commission is again tasked with drawing a plan. If the commission fails to enact a plan again, courts must step in and redraw district lines to ensure compliance with federal constitutional “one person, one vote” apportionment requirements. This is also considered a move by nature.
@@
All in all, the game specification depends on 19 parameters which govern the moves by nature, with most of these parameters relating to the court challenge process. Rather than fix these parameters to arbitrary constants, we place a prior distribution on each parameter over a range of probable values. We simulate 100 different draws from this joint prior distribution; each draw generates a slightly different game specification. We then average our results across the random draws.666We average, rather than propagating each draw through the estimation step. This has computational benefits and reduces the complexity of an outcome model (the outcome model would need to account for a different scale of treatment values across different draws). More importantly, however, the variation in the estimated equilibria, as described below, is quite small compared to other sources of uncertainty.
@@ 3.3 Equilibrium solution as a treatment variable
To effectively summarize a high-dimensional treatment, we use the subgame perfect Nash equilibrium of the redistricting game. The utility of the equilibrium solution captures the expected partisan bias of plans that arise out of a state’s redistricting process, under the current party control of the state’s institutions. All of the multi-step institutional interactions and negotiations that might happen as part of the redistricting process are thus reduced to a univariate score. The upshot is that the 14 procedural variables can be reduced into a single variable that measures the leeway political actors have over the partisan lean of the final redistricting plan.
[…]
The expected outcome of both the stalemate-then-court and court-alone choices are determined by the parameters governing the moves by nature. The overall equilibrium is a bias of , indicating that there is essentially no bias in favor of either party. This solution process is automated across all of the states and is carried out on each of the 100 different draws from the prior on the game’s parameters.
@@
For our causal analysis, the first treatment measure is exactly the game’s Nash equilibrium, averaged across the prior. This measure depends on which parties control each node in the game tree. We also generate a second treatment measure that does not depend on the current party control. This is calculated by assigning a single party (here, the Democrats777The game is completely symmetric between parties, with one exception: due to racial polarization in the U.S., challenges to redistricting plans under the VRA are much more likely for Republican-leaning plans, and when these challenges prevail the remedy almost always results in a plan more favorable to Democrats. In keeping with the goal of calculating the maximum leeway in each state, we therefore assign states to Democratic control under the counterfactual scenario here. Assigning to Republican control instead would slightly reduce the calculated leeway for those states previously subject to DOJ preclearance.) to control each node in the game tree that belongs to a partisan actor—legislature, governor, or partisan commission—and then recalculating the average Nash equilibrium. We refer to this equilibrium as the maximum leeway of a state’s redistricting process, since it captures the expected bias under a worst-case partisan outcome where all of the levers of state government are controlled by one party. The first treatment measure we refer to as the realized leeway, since it depends on the realized values of the party control for each state institution.
@@
As might be expected, states with independent redistricting commissions, such as the running example of Michigan, have a realized and maximum leeway of zero. In 2010, Michigan redistricting was entirely controlled by the legislature and governor, who were both Republican. By 2020, the process was reformed to have a strong independent commission. In the new process, the commission is bound by strict criteria, the process allows for court review, and even the commissioners themselves are selected in part by a lottery system. With the reform, Michigan’s maximum leeway changed from the highest observed value to zero, meaning that partisan actors have no control over redistricting outcomes, on average.
Other states’ scores are similarly interpretable, where introducing a commission (e.g., New York, Virginia, or Colorado) drastically reduces its leeway to match states with similar sets of rules in 2020. States that had intervening litigation to clarify the interpretation of state redistricting rules (e.g., New Mexico, North Carolina, or Pennsylvania) see appropriate changes in leeway, even absent a commission. Further, in our party-signed treatment, states that had a flip in party control of the whole system see large changes in leeway in the correct direction (e.g. West Virginia). Finally, states with minor changes to the total control of state institutions see minor changes in leeway (e.g., North Carolina).
@@ 3.4 Validation
[…]
The model was designed to be flexible enough to capture the actual redistricting processes in each state. Any misspecification is therefore due to how the moves by nature are specified, or in the underlying setup of two opposing partisan actors competing on a single zero-sum dimension. First, we find that the treatment values are not sensitive to specific parameter values, up to monotonic transformations. Across the 100 random draws from the prior, the average pairwise Spearman correlation between the Nash equilibria for each state is above 0.99.888i.e., for each pair of random draws, we calculate the Spearman correlation between the two vectors of calculated state equilibria, then average these correlations across all pairings. This also gives us confidence that the specific choice of the prior is not influencing the results.
Second, as shown in Figure 4, there is a substantial correlation in the expected direction between the treatment values and the measures of partisan advantage. The left panel plots the expected share of seats won by Republicans versus the realized leeway measure; recall that the leeway measure is positive for plans that favor Republicans. The right panel shows the partisan bias (King and Browning,, 1987), measured at the state’s baseline vote share, versus the realized leeway. The greater values of partisan bias correspond to plans that systematically favor one party, with positive values favoring Republicans.
For both outcome measures, the purely a priori model predictions of the expected partisan bias do in fact correlate with the actual partisan bias of the plans that come out of each state’s redistricting process. To what extent these correlations can tell us about the causal effects of reforms is of course another matter, and the primary question addressed in the next section.
@@
Table 1 compares the most likely final map-drawer, according to these probabilities, to the actual institution that drew the final redistricting plan. The correspondence is generally excellent, considering that the predictions are made based on theory alone. There is a tendency, however, for the model to predict court-drawn maps when they are in fact drawn by legislatures. The overall likelihood of a court intervention is itself a model parameter that is varied across random draws, and as noted above the ranking of the states by leeway is basically unchanged by different parameter values. Therefore, we expect this tendency to over-predict court intervention to at most slightly understate the equilibrium value for states with court review, since a higher likelihood of intervention would tend to pull the equilibrium towards zero.
Finally, it is natural to wonder whether a more accurate model could be obtained by estimating the model parameters from the observed data rather than specifying their prior distributions. That is, one could find the parameter values that maximize the correlation between the treatment and the outcome. We do not take this approach to cleanly separate the estimation of causal effects from the construction of a treatment variable (Rubin,, 2008). This enables us to employ a causal identification strategy that does not assume the correct specification of the game-theoretic model. Instead, we use the model to summarize a high-dimensional treatment variable in a theoretically informed way without looking at the outcome variable.999We did attempt to fit the game parameters to the observed data for the 2010 cycle alone via maximum likelihood, but we found that the observed data were only minimally informative about most of the parameter values. This is good news, insofar as it implies that a wide range of parameter values are all compatible with the observed data.
@@ 4.1 Difference-in-differences design
[…]
Our target estimand is the conditional average treatment effect (CATE) for a change from the treatment level to given a set of covariates that are not affected by the treatment. The CATE is defined as,
@@
Our covariates include a range of variables that existing work in the redistricting literature identifies as predicting redistricting outcomes and are not affected by redistricting reforms. We include (pre-treatment) 2008 Democratic presidential vote share, an indicator for being in the South,111111Southern states are defined as Alabama, Arkansas, Florida, Georgia, Kentucky, Louisiana, Mississippi, North Carolina, South Carolina, Tennessee, Texas, Virginia, and West Virginia. This set corresponds to Confederate states, plus Kentucky and West Virginia, which also had a history of Democratic dominance in the 20th century and notable differences between party preferences at the state and national levels. the logarithm of the number of districts in 2020, the change in the number of districts between 2010 and 2020, the logarithm of the average number of state corruption convictions by year between 2000 and 2010, and an indicator for whether states allow ballot initiatives. For the maximum leeway treatment, which uses no information about partisan control of state institutions, we also control for changes in partisan control of the map-drawing body and of the state supreme court.
Formally, we require the following conditional parallel trends assumption for all :
@@
4.2 Adjusting for political geography
To further increase the credibility of the conditional parallel trends assumption above, we adjust for the changes in political geography between the two redistricting cycles. Partisan preferences and the geographic distribution of voters can change in different ways within each state over time. For example, Michigan, which experienced reform, became more Republican during our sample period, while Georgia, which did not, became more Democratic. Without this adjustment, an estimated effect of the Michigan reform would be biased by the differential change in the states’ political geographies.